# LDM-ISP: Enhancing Neural ISP for Low Light with Latent Diffusion Models

### v3

@@ 4.5 Ablation Study
As we mentioned before, 2D discrete wavelet transforms (DWT) not only extract sub-bands for the taming but also downsample the input information to align with the latent size in the UNet. Instead of DWT, there are two other intuitive ways for downsampling: rescaling and CNN-based downsampling. In this study, we forego discussion on CNN-based downsampling due to its substantial computational expense. As shown in Fig. 4, simply resizing the input (e.g., using Bicubic Interpolation) results in the loss of many details and causes the small-scale input to be overwhelmed by noise. Taming the UNet with such input could result in an ambiguous global structure, potentially leading to low-quality sRGB images after decoding. A qualitative comparison between resizing and DWT taming strategies is shown in Fig. 7.
@@
The taming modules of the UNet hold significant importance as they steer the generating process toward generating specified latent representations. There is no need to conduct an ablation study on whether taming the UNet since the UNet without taming generates unexpected latent representation which is not related to the low-light input.
The fact that motivates us to tame the decoder $\mathcal{D}$ is that although the low-frequency sub-band (LL) helps the proposed method to tame the UNet for generating a high-correlation latent representation, the sRGB image decoded from this representation by a decoder lacking taming modules, shows much over-generated and irrelevant details.
Fig. 7 and Table 3 show that the decoder $\mathcal{D}$ when utilizing our DWT taming strategy, exhibits substantial improvements in color correction and high-frequency detail maintenance performance.
5 Conclusion

### v4

@@ IV-E Ablation Study
Our 2D discrete wavelet transforms (DWT) not only extract sub-bands for the taming but also downsample the input information to align with the latent size in the UNet. We also compare DWT with two other downsampling ways: rescaling and CNN-based downsampling. As shown in Fig. 2 (b), simply resizing the input (e.g., using Bicubic Interpolation) results in the loss of many details and causes the small-scale input to be overwhelmed by noise. Taming the UNet with such input could result in an ambiguous global structure, potentially leading to low-quality sRGB images after decoding. A qualitative comparison between resizing and DWT taming strategies is shown in Fig. 5.
@@
The taming modules of the UNet are crucial for guiding the generation process toward specific latent representations. An ablation study on taming is unnecessary since an untamed UNet produces irrelevant latent representations unrelated to low-light inputs.
We are motivated to tame the decoder $\mathcal{D}$ because, while the low-frequency sub-band (LL) aids in generating high-correlation latent representations, a decoder without taming modules results in over-generated and irrelevant details in the decoded sRGB image.
As shown in Fig.5 and Table II, using our DWT taming strategy significantly improves the decoder $\mathcal{D}$ in color correction and high-frequency detail maintenance.
V Conclusion
@@
VI Acknowledgement
This project was supported by the National Key R&D Program of China under grant number 2022ZD0161501.